(9/8)^-2 the 9/8 is of course a fraction and the ^-2 is for a negative exponent. I dont get these.
Answer by Alan3354(21612)

(9/8)^-2
----------
= (9/8)^-2/1
It's a fraction. Multiply NUM and DEN by (9/8)^2
= (9/8)^0/(9/8)^2
= 1/(9/8)^2 anything with exponent 0 = 1
--------------
Moving a term from NUM to DEN (or v.v.) changes the sign of the exponent.
(9/8)^-2 = 1/(9/8)^2
1/(9/8)^2 = (1/(9/8)^2 = (8/9)^2
1 over a fraction is the inverse of the fraction.
= 64/81